# Towards the use of multiple ROIs for radiomics-based survival modelling: finding a strategy of aggregating lesions

### v1

@@ 6 Conclusion
We proposed several methods of incorporating information from multiple lesions/regions of interest in radiomics-based survival models. We evaluated their performance with different types of models for the prediction of metastasis-free survival, based on PET images from a cohort of 115 non-small cell lung cancer patients. Regardless of the model, the predictive ability (assessed in terms of Harrell’s c-Index in 1000 validation iterations) could be improved by using aggregation of ROIs or risks, compared to a model built using only the primary tumour. Furthermore, including information from all lesions improved the robustness of the models.
Using all available feature vectors in a structurally heterogneous dataset to fit survival models can be considered a promising direction not only for imaging studies, but also for molecular research, in particular spatial or single-cell analysis.
@@ Acknowledgments
This work was supported by the Polish National Science Centre, grant number: UMO-2020/37/B/ST6/01959 (AS), Medical Research Agency, grant number: 07/010/ABB23/1037 (DB) and Silesian University of Technology statutory research funds (KF). AMW was also supported by a subsidy for the maintenance and development of research potential BKM. Calculations were performed on the Ziemowit computer cluster in the Laboratory of Bioinformatics and Computational Biology created in the EU Innovative Economy Programme POIG.02.01.00-00-166/08 and expanded in the POIG.02.03.01-00-040/13 project.

### v2

@@ 6 Conclusion
Incorporating other lesions in addition to the primary tumor, consistently improved the performance of survival models across various regression approaches. The aggregation strategies proposed in this work offer a robust solution for handling multiple radiomics ROIs, but also have broader applicability. Using all available feature vectors in a structurally heterogeneous dataset to fit survival models offers a promising direction, not only for imaging studies but also for molecular research, particularly in spatial or single-cell analysis.
@@ Acknowledgments
This work was supported by the Polish National Science Centre, grant number: UMO-2020/37/B/ST6/01959.Calculations were performed on the Ziemowit computer cluster in the Laboratory of Bioinformatics and Computational Biology created in the EU Innovative Economy Programme POIG.02.01.00-00-166/08 and expanded in the POIG.02.03.01-00-040/13 project.